Exercise 12-3 Internal Rate of Return [LO12-3]

Wendell’s Donut Shoppe is investigating the purchase of a new $50,100 donut-making machine. The new machine would permit the company to reduce the amount of part-time help needed, at a cost savings of $5,900 per year. In addition, the new machine would allow the company to produce one new style of donut, resulting in the sale of 2,600 dozen more donuts each year. The company realizes a contribution margin of $1.60 per dozen donuts sold. The new machine would have a six-year useful life.

Click here to view Exhibit 12B-1 and Exhibit 12B-2, to determine the appropriate discount factor(s) using tables.

Required:

1. What would be the total annual cash inflows associated with the new machine for capital budgeting purposes?

2. What discount factor should be used to compute the new machine’s internal rate of return? (Round your answer to 3 decimal places.)

3. What is the new machine’s internal rate of return? (Round your final answer to nearest whole percentage.)

4. In addition to the data given previously, assume that the machine will have a $17,225 salvage value at the end of six years. Under these conditions, what is the internal rate of return? (Hint: You may find it helpful to use the net present value approach; find the discount rate that will cause the net present value to be closest to zero.) (Round your final answer to nearest whole percentage.)

1. Annual cash inflows
2. Discount factor
3. Internal rate of return %
4. Internal rate of return %
Answer #1
1
Savings in part time help 5900
Additional contribution margin 4160 =2600*1.60
Annual cash inflows 10060
2
Discount factor 4.980 =50100/10060
3
The discount factor 4.980 for 6 years is closest to 6%
Internal rate of return 6%
4
Cash flows in year 4 27285 =10060+17225
Internal rate of return 12%